40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 3 04:01:55 2015, diego, Update, LSC, CM servo & AO path status CARM_03-02-2015_031754.pdf
    Reply  Tue Feb 3 16:36:33 2015, ericq, Update, LSC, CM servo & AO path status yarmAO.pdf
       Reply  Mon Feb 9 17:47:17 2015, diego, Update, LSC, CM servo & AO path status 
Message ID: 10966     Entry time: Tue Feb 3 04:01:55 2015     Reply to this: 10969
Author: diego 
Type: Update 
Category: LSC 
Subject: CM servo & AO path status 

[Diego, Jenne]

Tonight we worked on the CM board and AO path:

  • at first we changed the REFL1 input to the CM board from REFL11 to AS55, as written in my previous elog; we tried following Koji's procedures from http://nodus.ligo.caltech.edu:8080/40m/9500 but we didn't get any result: we could lock using the regular digital path but no luck at all for the analog path;
  • then we decided to follow the procedure to the letter, using POY11Q as input to the CM board;
    • we still couldn't lock following the Path #2, even after adjusting the gains to match the current configuration for the Yarm filter bank;
    • we had some more success using Path #1, but we had to lower the REFL1 Gain to ~3-4 (from the original 31) because of the different configuration of the Yarm filter bank, in order to have the same sensing in both of them; we managed to acquire lock a few times, it's not super stable but it can keep lock for a while;
    • when we tried to increase the gain of the MC filter bank and the AO Gain, however, we immediately had some gain peaking, and we couldn't go further then 0.15 and 9db respectively. We currently don't have an answer for that.
    • anyhow, we took a few measurements with the SR785:

 

The BLUE plot is at MC Gain = 0.10 and REFL1 Gain = 4dB; the GREEN plot is for MC Gain = 0.10 and REFL1 Gain = 3dB, which seemed a more stable configuration; after this last configuration, we increased the MC Gain to 0.15 and the AO Gain from 8dB to 9dB and took another measurement, the RED plot; this is as far as we got as of now. We also couldn't increase the REFL11 Gain because it made things unstable and more prone to unlock.

So, some little progress on the AO path procedure, but we are very low on our UGF and we have to find a way to increase our gains without breaking the lock and avoiding the gain peaking we have witnessed tonight.

 

Notes:

  • is the REFL1 Gain dB slider supposed to go to negative dBs? During the night we also tried to use negative dBs, but it seemed it wasn't doing anything instead;
  • when we plugged POY11Q to the CM board, we noticed that it wasn't connected to anything at the moment; since we phase rotate POY11, we were assuming that we were using that signal somewhere. We are confused by this...
  • we remind that REFL11 is no more connected to the CM board input, as POY11 is.
Attachment 1: CARM_03-02-2015_031754.pdf  35 kB  | Hide | Hide all
CARM_03-02-2015_031754.pdf
ELOG V3.1.3-